# Fluctuations of the Ising free energy on Erdős-Rényi graphs

### v1

@@ 2.3. The central limit theorem in the case $B=0$
Part (iii) of the proposition, which clarifies the combinatorial meaning of the function $r_{d,\beta}$, is an easy consequence of statements from [11, 23].
@@
We emphasise that the conditioning on $\mathcal{S}_{+}$ is crucial in Proposition 2.15. Due to inversion symmetry the unconditional magnetisations just satisfy $\left\langle{{\boldsymbol{\sigma}(v_{i})},{\mu_{\mathbb{G},\beta,0}}}\right\rangle=0$; thus, they are uninformative. By contrast, Proposition 2.14 (ii) and Proposition 2.15 (iii) imply that the empirical distribution $\boldsymbol{\pi}^{+}_{n,m,\beta}$ of the conditional magnetisations converges to a limiting distributions with a strictly positive mean. This fact together with the replica symmetry provided by Proposition 2.15 (ii) easily implies that for the unconditional measure $\mu_{\mathbb{G},\beta,0}$ we have
@@
In other words, the unconditional $\mu_{\mathbb{G},\beta,0}$ genuinely fails to be replica symmetric. The proof of Proposition 2.15 combines the insights from Proposition 2.14 and techniques for the study of diluted spin glass models [7, 18, 19, 20] with somewhat delicate monotonicity and positive correlation arguments.
@@ 2.4. The case $B=0$, high temperature
Theorem 1.2 asserts that in the high temperature scenario without an external field, i.e., $B=0$ and $d\leq 1$ or $\beta<\beta_{\mathrm{f}}(d)$ the free energy only exhibits bounded fluctuations. To prove this fact we employ the small subgraph conditioning method [36, 57]. Thus, we are going to argue that the variance of the partition function $Z_{\mathbb{G}}(\beta,0)$ can be attributed entirely to the presence of bounded-length cycles in $\mathbb{G}$. Since it is well known that the number of cycles of a fixed length $\ell\geq 3$ in $\mathbb{G}$ converges to $\boldsymbol{c}_{\ell}\sim{\rm Po}(d^{\ell}/(2\ell))$, we will thus obtain the limiting distribution stated in Theorem 1.2. The proofs for the following statements can be found in Section 6, where we prove Propositions 2.19, 2.20 and 2.21, from which we ultimately derive Theorem 1.2.
@@ 2.5. Discussion
As pointed out at the start of the section, the contribution most closely related to the present work on a technical level is the recent article by Chatterjee, Coja-Oghlan, Müller, Riddlesden, Rolvien, Zakharov and Zhu on the random 2-SAT problem [14]. Specifically, that article shows that throughout the satisfiable phase of the random 2-SAT problem the number of satisfying assignments is asymptotically log-normal. Similarly to the present work (see the coupling CLP0–CPL2), the proof is based on investigating pairs of correlated random 2-SAT instances that share a certain number of clauses and each contain a number of independent random clauses. That proof strategy, in turn, is inspired by the proof of the central limit theorem for fully connected spin glass models by Chen, Dey and Panchenko [15], which is based on an interpolating family of Hamiltonians.
@@
By contrast, in the case $B=0$ matters get more involved. For a start, for $d>1$ and $\beta>\beta_{\mathrm{f}}(d)$ it is not just that the Gibbs measure on the Galton-Watson tree is no longer unique, but the free boundary condition renders an uninformative and unstable Belief Propagation fixed point distribution (the atom at zero). Furthermore, as we pointed out in Eq. (2.19), the replica symmetry condition (2.3) does not hold in this case. Instead, we need to derive the pure state decomposition as stated in Proposition 2.15. The proof of this proposition constitutes the main technical challenge of the present work. To elaborate, the proof of Proposition 2.15 relies on a general information theoretic decomposition result for probability measures called the ‘pinning lemma’ from [18], see Lemma 3.9 below. The pinning lemma is a generalisation of earlier results from Montanari [51] and Raghavendra and Tan [56]. It was seized upon in prior work [19, 20] to study diluted models of disordered systems such as spin glass models. The pinning lemma implies that the Ising model on $\mathbb{G}$ admits a ‘simple’ decomposition into a ‘small’ number of pure states where the conditional Ising distribution enjoys replica symmetry. Starting from the pinning lemma, what we are going to show by means of monotonicity and positive correlation arguments is that the only pure state decomposition that can be reconciled with the free energy formula (1.5) is the one stated in Proposition 2.15. Once this is in place we will then show that in the proof of the central limit theorem the pure state decomposition can be used in lieu of the Gibbs uniqueness property.
@@ 3.4. Small subgraph conditioning
The proof of Theorem 1.2 is based the small subgraph conditioning technique for which the following theorem provides the general framework.
@@
Now, pick $\delta=\delta(\varepsilon,\ell)>0$ small enough and suppose that $t^{\prime},t^{\prime\prime}\in[0,1]$ satisfy $|t^{\prime}-t^{\prime\prime}|<\delta$. Then for either value $t\in\{t^{\prime},t^{\prime\prime}\}$ the definition of the multi-type Galton-Watson tree $\mathbb{T}^{\otimes}$ provides that a type 0 vertex spawns ${\rm Po}(d(1-t))$ offspring of type 0, and ${\rm Po}(dt)$ offspring of type 1 and type 2. (The offspring distributions of type 1 and type 2 vertices are independent of $t$.) Hence, providing that $|t^{\prime}-t^{\prime\prime}|<\delta$ for a sufficiently small $\delta=\delta(\varepsilon,\ell)>0$, we can couple these Poisson variables such that
@@
Thus, we obtain GeG-e from $G$ by removing the edge $e$, then impose a +1+1 ‘boundary condition’ on all vertices $u$ at distance precisely $\ell$ form $v$ in GeG-e and finally evaluate the magnetisation of $v$. The definition (5.2) is similar to the idea of a ‘standard Belief Propagation message’ [20, 48]. We obtain the following bound.
@@
Moreover, again because the distance between $v,w$ in GeG-e exceeds $2\ell$, the spins $\boldsymbol{\sigma}(v),\boldsymbol{\sigma}(w)$ are stochastically independent under once we impose a +1+1 boundary condition on $\partial^{\ell}(G-e,v)\cup\partial^{\ell}(G-e,w)$. Therefore,
@@
Assume that $d>1$ and $\beta\in(\beta_{\mathrm{f}}(d),\infty)\setminus\mathfrak{X}_{d}$. Then for any fixed $\delta,\ell>0$ the following is true the event $\mathfrak{D}_{\delta,\ell}$ that
@@
Furthermore, providing $\xi$ is sufficiently small, Lemmas 3.6 and 3.7 show that the measure $\mu_{\mathbb{G},\beta,0}(\,\cdot\,\mid\mathcal{S}_{\tau})$ is $\zeta$-symmetric. Hence, (5.42) yields
@@
Combining (6.4) and (6.5), we obtain the expression for (6.1). To derivation of (6.2) is similar, except that we only sum on configurations $\sigma$ with $|\sum_{v\in V_{n}}\sigma(v)|\leq 2\sqrt{n}\log n$. ∎
@@
$=\frac{\Theta(n^{-1/2})}{\exp(\beta m)}\sum_{m_{+}+m_{-}\leq m}\binom{m}{m_{+},m_{-}}\left({\frac{|\sigma^{-1}(1)|}{n}}\right)^{2m_{+}}\left({\frac{|\sigma^{-1}(-1)|}{n}}\right)^{2m_{-}}\left({\frac{2|\sigma^{-1}(1)|\cdot|\sigma^{-1}(-1)|}{n^{2}}}\right)^{m-m_{+}-m_{-}}\exp(2\beta(m_{+}+m_{-}))$

### v2

@@ 2.3. The central limit theorem in the case $B=0$
Part (iii) of the proposition, which clarifies the combinatorial meaning of the function $r_{d,\beta}$, is an easy consequence of statements from [11, 23]; see Section 5.1 for details.
@@
We emphasise that the conditioning on $\mathcal{S}_{+}$ is crucial in Proposition 2.15. Due to inversion symmetry the unconditional magnetisations just satisfy $\left\langle{{\boldsymbol{\sigma}(v_{i})},{\mu_{\mathbb{G},\beta,0}}}\right\rangle=0$; thus, they are uninformative. By contrast, Proposition 2.14 (ii) and Proposition 2.15 (iii) imply that the empirical distribution $\boldsymbol{\pi}^{+}_{n,m,\beta}$ of the conditional magnetisations converges to a limiting distribution with a strictly positive mean. This fact together with the replica symmetry provided by Proposition 2.15 (ii) easily implies that for the unconditional measure $\mu_{\mathbb{G},\beta,0}$ we have
@@
In other words, the unconditional $\mu_{\mathbb{G},\beta,0}$ genuinely fails to be replica symmetric. The proof of Proposition 2.15 combines the insights from Proposition 2.14 and techniques for the study of diluted spin glass models [7, 18, 19, 20] with monotonicity and positive correlation arguments.
@@ 2.4. The case $B=0$, high temperature
Theorem 1.2 asserts that in the high temperature scenario without an external field, i.e., $B=0$ and $d\leq 1$ or $\beta<\beta_{\mathrm{f}}(d)$ the free energy only exhibits bounded fluctuations. To prove this fact we employ the small subgraph conditioning method [36, 57]. Thus, we are going to argue that the variance of the partition function $Z_{\mathbb{G}}(\beta,0)$ can be attributed entirely to the presence of bounded-length cycles in $\mathbb{G}$. Since it is well known that the number of cycles of a fixed length $\ell\geq 3$ in $\mathbb{G}$ converges to $\boldsymbol{c}_{\ell}\sim{\rm Po}(d^{\ell}/(2\ell))$, we will thus obtain the limiting distribution stated in Theorem 1.2. The proofs for the following statements can be found in Section 6, where we prove Propositions 2.19, 2.20 and 2.21.
@@ 2.5. Discussion
As pointed out at the start of the section, the contribution most closely related to the present work on a technical level is the recent article by Chatterjee, Coja-Oghlan, Müller, Riddlesden, Rolvien, Zakharov and Zhu on the random 2-SAT problem [14]. Specifically, that article shows that throughout the satisfiable phase of the random 2-SAT problem the number of satisfying assignments is asymptotically log-normal. Similarly to the present work (see the coupling CPL0–CPL2), the proof is based on investigating pairs of correlated random 2-SAT instances that share a certain number of clauses and each contain a number of independent random clauses. That proof strategy, in turn, is inspired by the proof of the central limit theorem for fully connected spin glass models by Chen, Dey and Panchenko [15], which is based on an interpolating family of Hamiltonians.
@@
By contrast, in the case $B=0$ matters get more involved. For a start, for $d>1$ and $\beta>\beta_{\mathrm{f}}(d)$ it is not just that the Gibbs measure on the Galton-Watson tree is no longer unique, but the free boundary condition renders an uninformative and unstable Belief Propagation fixed point distribution (the atom at zero). Furthermore, as we pointed out in Eq. (2.19), the replica symmetry condition (2.3) does not hold in this case. Instead, we need to derive the pure state decomposition as stated in Proposition 2.15. The proof of this proposition constitutes the main technical challenge of the present work. To elaborate, the proof of Proposition 2.15 relies on a general information theoretic decomposition result for probability measures called the ‘pinning lemma’ from [18], see Lemma 3.9 below. The pinning lemma is a generalisation of earlier results from Montanari [51] and Raghavendra and Tan [56]. It was seized upon in prior work [19, 20] to study diluted models of disordered systems such as spin glass models. The pinning lemma implies that the Ising model on $\mathbb{G}$ admits a ‘simple’ decomposition into a ‘small’ number of pure states where the conditional Ising distribution enjoys replica symmetry. Starting from the pinning lemma, ultimately we are going to argue by means of monotonicity and positive correlation arguments that the only pure state decomposition that can be reconciled with the free energy formula (1.5) is the one stated in Proposition 2.15. Once this is in place we will then show that in the proof of the central limit theorem the pure state decomposition can be used in lieu of the Gibbs uniqueness property.
@@ 3.4. Small subgraph conditioning
The proof of Theorem 1.2 is based on the small subgraph conditioning technique for which the following theorem provides the general framework.
@@
Now, pick $\delta=\delta(\varepsilon,\ell)>0$ small enough and suppose that $t^{\prime},t^{\prime\prime}\in[0,1]$ satisfy $|t^{\prime}-t^{\prime\prime}|<\delta$. Then for either value $t\in\{t^{\prime},t^{\prime\prime}\}$ the definition of the multi-type Galton-Watson tree $\mathbb{T}^{\otimes}$ provides that a type 0 vertex spawns ${\rm Po}(d(1-t))$ offspring of type 0, and ${\rm Po}(dt)$ offspring of type 1 and type 2. (The offspring distributions of type 1 and type 2 vertices are independent of $t$.) Hence, provided that $|t^{\prime}-t^{\prime\prime}|<\delta$ for a sufficiently small $\delta=\delta(\varepsilon,\ell)>0$, we can couple these Poisson variables such that
@@
Thus, we obtain GeG-e from $G$ by removing the edge $e$, then impose a +1+1 ‘boundary condition’ on all vertices $u$ at distance precisely $\ell$ from $v$ in GeG-e and finally evaluate the magnetisation of $v$. The definition (5.2) is similar to the idea of a ‘standard Belief Propagation message’ [20, 48]. We obtain the following bound.
@@
Moreover, again because the distance between $v,w$ in GeG-e exceeds $2\ell$, the spins $\boldsymbol{\sigma}(v),\boldsymbol{\sigma}(w)$ are stochastically independent once we impose a +1+1 boundary condition on $\partial^{\ell}(G-e,v)\cup\partial^{\ell}(G-e,w)$. Therefore,
@@
Assume that $d>1$ and $\beta\in(\beta_{\mathrm{f}}(d),\infty)\setminus\mathfrak{X}_{d}$. Then for any fixed $\delta,\ell>0$ the event $\mathfrak{D}_{\delta,\ell}$ that
@@
Furthermore, provided $\xi$ is sufficiently small, Lemmas 3.6 and 3.7 show that the measure $\mu_{\mathbb{G},\beta,0}(\,\cdot\,\mid\mathcal{S}_{\tau})$ is $\zeta$-symmetric. Hence, (5.42) yields
@@
Combining (6.4) and (6.5), we obtain the expression for (6.1). The derivation of (6.2) is similar, except that we only sum on configurations $\sigma$ with $|\sum_{v\in V_{n}}\sigma(v)|\leq 2\sqrt{n}\log n$. ∎
@@
$=\frac{\Theta(n^{-1/2})}{\exp(\beta m)}\sum_{m_{+}+m_{-}\leq m}\binom{m}{m_{+},m_{-}}\left({\frac{|\sigma^{-1}(1)|}{n}}\right)^{2m_{+}}\left({\frac{|\sigma^{-1}(-1)|}{n}}\right)^{2m_{-}}$
$\qquad\qquad\qquad\qquad\quad\quad\cdot\left({\frac{2|\sigma^{-1}(1)|\cdot|\sigma^{-1}(-1)|}{n^{2}}}\right)^{m-m_{+}-m_{-}}\exp(2\beta(m_{+}+m_{-}))$